# Using Pre-training and Interaction Modeling for ancestry-specific disease prediction using multiomics data from the UK Biobank Thomas Le Menestrel1*, Erin Craig3, Robert Tibshirani2,3, Trevor Hastie2, Manuel Rivas3*

### v1

@@ Abstract
Recent genome-wide association studies (GWAS) have uncovered the genetic basis of complex traits, but show an under-representation of non-European descent individuals, underscoring a critical gap in genetic research. Here, we assess whether we can improve disease prediction across diverse ancestries using multiomic data. We evaluate the performance of Group-LASSO INTERaction-NET (glinternet) and pretrained lasso in disease prediction focusing on diverse ancestries in the UK Biobank. Models were trained on data from White British and other ancestries and validated across a cohort of over 96,000 individuals for 8 diseases. Out of 96 models trained, we report 16 with statistically significant incremental predictive performance in terms of ROC-AUC scores (). These findings suggest that advanced statistical methods that borrow information across multiple ancestries may improve disease risk prediction, but with limited benefit. Our code is available at (https://github.com/rivas-lab/AncestryOmicsUKB).
@@
Glinternet and pretrained lasso are statistical methods that excel in scenarios where data for one population is limited, by leveraging common patterns from a larger population through interaction modelling and pre-training. This capability is crucial for the UK Biobank where 75% of the data pertains to White British individuals, and less than 10% to each non-White British ancestry. Despite the widespread use of LASSO and Cox proportional hazard models, strategies that consider interactions and pre-training remain unexplored [1, 2]. In this study, we trained glinternet and pretrained lasso models on mixed datasets of White British and other ancestries including South Asian, African, non-British European and Admixed across 8 diseases, reporting 16 models with significant improvements () out of 96. Our findings indicate that both interaction terms and pre-training can enhance prediction accuracy but for a limited set of diseases.
@@ Introduction
Genome-wide association studies (GWAS) have significantly advanced our understanding of the genetic basis of complex traits and diseases [3, 4]. Yet, these studies have predominantly focused on individuals of European descent, leading to under-representation of non-European ancestry in genomic databases [5, 6], despite ancestry being a crucial factor in disease susceptibility [7]. This is also the case for the UK Biobank, which contains approximately 75% data from White British individuals and less than 5% from diverse ancestries such as South Asian, African, and East Asian [8].
@@
Efforts to address the challenge of ancestry-specific disease prediction have varied. Polygenic risk scores (PRS) were trained for disease prediction using UK Biobank data [9, 10]; however, their efficacy diminishes when applied to non-European individuals. Alternative strategies employing LASSO variants and advanced machine learning techniques like AutoEncoders have been attempted [11, 12]. Yet, these methods struggle with non-linearity and lack the interpretability of simpler models, limiting their utility. Cox proportional hazard models, as well as penalized variants, have been extensively used for this problem but suffer from the same issues as LASSO models when dealing with non linearity as well as from the proportional hazard assumption failing in practice [13, 14, 15]. Finally, how multiomics data beyond genetics improve disease risk prediction in non-European populations is unclear and most studies have focused on single data entity [16, 17, 18].
@@
Addressing these limitations, our study uses pre-training and interaction modeling through glinternet and pretrained lasso across multiple ancestries and individual ancestries jointly, and facilitate the application of patterns identified in White British data to other ancestries. This approach not only improves prediction accuracy across a diverse range of ancestries but also maintains the level of interpretability achieved in previous studies. We compiled a dataset of over 93,000 individuals spanning 8 diseases with combined demographic, metabolomic, genetic, and biomarker data.
-0.2in0in
Ancestry | Cases | DIA | MI | AST | GS | OST | ART | CYS | CRF |
WB | Positive | 5479 | 3234 | 10909 | 4103 | 7579 | 4969 | 2725 | 2862 |
WB | Negative | 75331 | 77576 | 69901 | 76707 | 73231 | 75841 | 78085 | 77948 |
WB | Total | 80810 | 80810 | 80810 | 80810 | 80810 | 80810 | 80810 | 80810 |
SA | Positive | 460 | 151 | 285 | 69 | 101 | 122 | 56 | 93 |
SA | Negative | 1451 | 1760 | 1626 | 1842 | 1810 | 1789 | 1855 | 1818 |
SA | Total | 1911 | 1911 | 1911 | 1911 | 1911 | 1911 | 1911 | 1911 |
AF | Positive | 232 | 30 | 201 | 36 | 86 | 103 | 24 | 83 |
AF | Negative | 1267 | 1469 | 1298 | 1463 | 1413 | 1396 | 1475 | 1416 |
AF | Total | 1499 | 1499 | 1499 | 1499 | 1499 | 1499 | 1499 | 1499 |
AD | Positive | 632 | 268 | 932 | 285 | 492 | 379 | 234 | 213 |
AD | Negative | 6151 | 6515 | 5851 | 6498 | 6291 | 6404 | 6549 | 6570 |
AD | Total | 6783 | 6783 | 6783 | 6783 | 6783 | 6783 | 6783 | 6783 |
NBE | Positive | 419 | 197 | 768 | 262 | 433 | 279 | 178 | 165 |
NBE | Negative | 5491 | 5713 | 5142 | 5648 | 5477 | 5631 | 5732 | 5745 |
NBE | Total | 5910 | 5910 | 5910 | 5910 | 5910 | 5910 | 5910 | 5910 |
@@
Ethics statement
This research has been conducted using the UK Biobank Resource under Application Number 24983, βGenerating effective therapeutic hypotheses from genomic and hospital linkage dataβ (http://www.ukbiobank.ac.uk/wp-content/uploads/2017/06/24983-Dr-Manuel-Rivas.pdf).
Based on the information provided in Protocol 44532, the Stanford IRB has determined that the research does not involve human subjects as defined in 45 CFR 46.102(f) or 21 CFR 50.3 (g). All participants of the UK Biobank provided written informed consent (information available at https://www.ukbiobank.ac.uk/2018/02/gdpr/).
@@
Pretrained lasso
In Generalized Linear Models (GLMs) and -regularized GLMs, an offsetβa predetermined -vectorβcan be added as an extra column in the feature matrix with a fixed weight of 1 and allows to control for a variable without estimating its parameter [19]. Additionally, the standard norm can be extended to a weighted norm by assigning a penalty factor to each feature, modulating the regularization applied [20]. A penalty factor of zero indicates no penalty, ensuring feature inclusion, while infinity results in feature exclusion, enhancing model flexibility and specificity.
@@
The UK Biobank is a comprehensive cohort study that gathers data from various locations across the UK. To address the potential variability caused by population structure within our dataset, we limited our analysis to individuals who are not related, adhering to four specific criteria outlined in the UK Biobankβs sample quality control file, ukb_sqc_v2.txt. These criteria are: (1) inclusion in principal component analysis (as indicated in the used_in_pca_calculation column); (2) individuals not identified as outliers based on heterozygosity and missing data rates (found in the het_missing_outliers column); (3) absence of suspected anomalies in sex chromosome number (mentioned in the putative_sex_chromosome_aneuploidy column); and (4) having no more than ten likely third-degree relatives (as per the excess_relatives column).
We refined our study population by analyzing genetic and self-reported ancestry data, using the UK Biobankβs genotype principal components (PCs), self-reported ancestry data (UK Biobank Field ID 21000), and the in_white_British_ancestry_subset column from the sample QC file. Our focus was on individuals self-identifying as white British (337,129), non-British European (24,905), African (6,497), South Asian (7,831), and East Asian (1,704). The classification into five groups involved a two-stage process, initially using genotype PC loadings to apply specific criteria for each group: (1) white British based on PC1 and PC2 thresholds and inclusion in the white British ancestry subset; (2) non-British European with similar PC thresholds, identified as white but not white British; (3) African, with distinct PC thresholds, excluding identities as Asian, White, Mixed, or Other; (4) South Asian and (5) East Asian, each with unique PC thresholds and exclusions of identifying as Black, White, Mixed, or Other.
@@
Our cohort consisted of 80,810 White British, 1,911 South Asians, 1,499 Africans, 6,783 categorized as Admixed, and 5,910 Non-British Europeans (see Table 2 for detailed breakdown). We identified 8 common diseases to use for our study for which we had a minimum of 20 positive cases per ancestry: diabetes, asthma, chronic renal failure, osteoarthritis, myocardial infarction, cystistis, gall stones and arthritis.
@@
The dataset labeled βAllβ contains data from all ancestries available in the UK Biobank, including White British (WB) as well as ancestries South Asian, African, Admixed and Non-British European. The βMixβ dataset refers to a combination of White British data with data from one other specific ancestry at a time (e.g. White British and South Asian). This approach allows us to explore the predictive performance of our models when they are trained on data representing the majority population (White British) alongside data from a minority ancestry. The term βancestry-only dataβ is used to denote datasets consisting exclusively of data from a single, specific ancestryβsuch as South Asian (SA) for instance. By employing these distinct dataset configurations, our study aims to assess whether training models on all ancestries is beneficial compared to training on White British data and data from a single ancestry.
@@
We started our analysis with a baseline model β an L1-penalized logistic regression β identified as the standard approach for this type of analysis in our literature review [2]. We used the glmnet package [21] in R [22] and trained it with the three distinct datasets aforementioned: ancestry-only data, mixed data, and data encompassing all ancestries (see Figure 1).
Subsequently, we applied the glinternet package in R to train glinternet models [23], selecting age, sex, ancestry, and PRS score as interaction candidates (see Figure 1). Training of glinternet models is conducted on both the Mix and All datasets to assess glinternetβs ability to extrapolate patterns across ancestries facing data limitations.
Last, we adopted a two-stage process to train the pretrained lasso model. We used the ptLasso package in R [24] to train a generic LASSO model using glmnet, from which we derive the vector. This vector is then utilized to construct an offset parameter, facilitating the training of ancestry-specific models. A comprehensive pan-ancestry model is subsequently developed by combining these steps (see Figure 1) and is trained like the glinternet models on both the Mix and All datasets.

### v2

@@ Abstract
Recent genome-wide association studies (GWAS) have uncovered the genetic basis of complex traits, but show an under-representation of non-European descent individuals, underscoring a critical gap in genetic research. Here, we assess whether we can improve disease prediction across diverse ancestries using multiomic data. We evaluate the performance of Group-LASSO INTERaction-NET (glinternet) and pretrained lasso in disease prediction focusing on diverse ancestries in the UK Biobank. Models were trained on data from White British and other ancestries and validated across a cohort of over 96,000 individuals for 8 diseases. Out of 96 models trained, we report 16 with statistically significant incremental predictive performance in terms of ROC-AUC scores (), found for diabetes, arthritis, gall stones, cystitis, asthma and osteoarthritis. For the interaction and pretrained models that outperformed the baseline, the PRS score was the primary driver behind prediction. Our findings indicate that both interaction terms and pre-training can enhance prediction accuracy but for a limited set of diseases and moderate improvements in accuracy. Our code is available at (https://github.com/rivas-lab/AncestryOmicsUKB).
@@
Glinternet and pretrained lasso are statistical methods that excel in scenarios where data for one population is limited, by leveraging common patterns from a larger population through interaction modelling and pre-training. This capability is crucial for the UK Biobank where 75% of the data pertains to White British individuals while certain ancestries make up less than 1% of the database. Despite the widespread use of LASSO and Cox proportional hazard models [1, 2], strategies that consider interactions and pre-training remain unexplored. In this study, we trained glinternet and pretrained lasso models on mixed datasets of White British and other ancestries including South Asian, African, non-British European and Admixed across 8 diseases, reporting 16 models with significant improvements () out of 96.
@@ Introduction
Genome-wide association studies (GWAS) have significantly advanced our understanding of the genetic basis of complex traits and diseases [3, 4]. Yet, these studies have predominantly focused on individuals of European descent, leading to under-representation of non-European ancestry in genomic databases [5, 6], despite ancestry being a crucial factor in disease susceptibility [7]. This is also the case for the UK Biobank, which contains approximately 75% data from White British individuals and less than 10% from diverse ancestries such as Related (9.9%), Admixed (6.4%), Non-British European (5.6%), South Asian (1.8%), African (1.5%) and East Asian (0.4%) [8].
@@
Efforts to address the challenge of ancestry-specific disease prediction have varied. Polygenic risk scores (PRS) were trained for disease prediction using UK Biobank data [9, 10]; however, their efficacy diminishes when applied to non-European individuals. PCA approaches have also been developed to create ancestry-specific PRS scores for disease prediction [11]. Alternative strategies employing LASSO variants and advanced machine learning techniques like AutoEncoders have been attempted [12, 13]. Yet, these methods struggle with non-linearity and lack the interpretability of simpler models. LASSO and Cox proportional hazard models, as well as penalized variants, have been extensively used for this problem but under-perform when dealing with non-linearity as well as because of the proportional hazard assumption failing in practice [14, 15, 16, 17]. Finally, how multiomics data beyond genetics improve disease risk prediction in non-European populations is unclear and most studies have focused on single data entity as well as single diseases [18, 19, 20].
Addressing these limitations, our study uses pre-training and interaction modeling through glinternet and pretrained lasso across multiple ancestries and individual ancestries jointly, and facilitate the application of patterns identified in White British data to other ancestries. Interaction modeling allows to adjust coefficients of certain variables based on specific covariates, often age, sex or PRS score for a given disease. This allows to capture non-linear relationships between features, for instance, the effect of a genetic variant which might be different depending on the ancestry of a patient or environmental factors. On the other hand, pre-training on a large dataset with multiple ancestries is particularly beneficial when underlying genetic patterns may be complex and not easily discernible from smaller datasets. After the initial broad training, the model is fine-tuned on a smaller, more specific dataset. This is crucial when dealing with PRS scores for diverse ancestries because it allows the model to adapt to the nuances of specific genetic backgrounds that might not be well-represented in the larger dataset.
@@
We compiled clinical data from over 96,000 individuals spanning 8 diseases with demographic, metabolomic, genetic and biomarker data, thus having a dataset with multiomic data and multiple diseases, addressing the issues seen in previous studies. This approach improves prediction accuracy across a diverse range of ancestries and maintains the level of interpretability achieved in previous studies.
Ethics statement
This research has been conducted using the UK Biobank Resource under Application Number 24983, “Generating effective therapeutic hypotheses from genomic and hospital linkage data” (http://www.ukbiobank.ac.uk/wp-content/uploads/2017/06/24983-Dr-Manuel-Rivas.pdf). Based on the information provided in Protocol 44532, the Stanford IRB has determined that the research does not involve human subjects as defined in 45 CFR 46.102(f) or 21 CFR 50.3 (g). All participants of the UK Biobank provided written informed consent (information available at https://www.ukbiobank.ac.uk/2018/02/gdpr/).
Results
Results for African ancestry
For African ancestry, we find 6 statistically significant models (), which are all glinternet models (cystitis, gall stones, diabetes and arthritis). In terms of average ROC-AUC score across all diseases, pretrained lasso shows suboptimal performance and is similar or worse than the baseline models depending on the subset used for training, while glinternet slighty outperforms both the logistic regression and pretrained lasso models (Tables 2 and 3).
The glinternet model shows significant improvements over the baseline for several diseases such as cystitis ( ROC-AUC of 0.039, 7.1% increase), gall stones ( ROC-AUC of 0.054, 11.2 % increase), diabetes ( ROC-AUC of 0.022, 2.5 % increase) and arthritis ( ROC-AUC of 0.073, 11.4 % increase). The model is consistent across all diseases and is only slighty outperformed for myocardial infarction and osteoarthritis by a small margin (less than 0.01 ROC-AUC). The glinternet models trained on the Mix datasets have an average ROC-AUC 0.028 higher than the baselines (4.1 % increase).
We find that L1-penalized Logistic Regression keeps 147 non-zero coefficients when trained on the Mix dataset for diabetes, while Pretrained LASSO keeps only 19 of them and outperforms it by a ROC-AUC of 0.02. For Glinternet for the same disease, we report a ROC-AUC of 0.022 and find 60 interaction terms (27 with age, 25 with the PRS score for diabetes and 8 with sex).
Glinternet and logistic regression provide the best accuracy for predicting certain diseases in individuals of African ancestry, specifically for cystitis, gall stones, diabetes and arthritis, as shown by the ROC-AUC scores and the p-values of the ROC tests. On the contrary, pretrained lasso does not show statistically significant improvements for the ROC tests and has same or lower average ROC-AUC scores than both baselines (Tables 1 and 2).
-1.85in0in Method Data CRF CYS GL DIA MI OST AST ART Avg. Inc. Logistic regression AF 0.718 0.5 0.5 0.844 0.5 0.797 0.622 0.718 0.65 0 Logistic regression Mix 0.71 0.549 0.474 0.839 0.753 0.847 0.629 0.641 0.68 0.03 Logistic regression All 0.705 0.549 0.471 0.854 0.755 0.848 0.643 0.683 0.688 0.038 Glinternet Mix 0.706 0.588 0.527 0.86 0.754 0.84 0.675 0.714 0.708 0.058 Glinternet All 0.718 0.526 0.501 0.872 0.745 0.839 0.667 0.741 0.701 0.051 Pretrained lasso Mix 0.689 0.568 0.478 0.858 0.63 0.788 0.639 0.697 0.669 0.018 Pretrained lasso All 0.725 0.567 0.477 0.865 0.646 0.838 0.621 0.701 0.68 0.03
-1.25in0in Method Metric Data CRF CYS GL DIA MI OST AST ART Glinternet P-value Mix 0.572 0.016 0.033 0.002 0.470 0.704 0.073 0.018 Glinternet AUC Mix -0.004 0.039 0.054 0.022 0.001 -0.008 0.046 0.073 Glinternet P-value All 0.250 0.653 0.079 0.025 0.729 0.728 0.166 0.012 Glinternet AUC All 0.013 -0.023 0.030 0.018 -0.010 -0.009 0.024 0.058 Pretrained lasso P-value Mix 0.655 0.463 0.460 0.073 0.844 0.920 0.418 0.101 Pretrained lasso AUC Mix -0.021 0.019 0.005 0.02 -0.123 -0.059 0.01 0.056 Pretrained lasso P-value All 0.274 0.221 0.440 0.138 0.948 0.735 0.665 0.176 Pretrained lasso AUC All 0.02 0.018 0.006 0.01 -0.109 -0.009 -0.022 0.018
Results for South Asian ancestry
For South Asian ancestry, we find 8 statistically significant models (): five glinternet models (osteoarthritis, arthritis, asthma) and 3 pretrained lasso models (asthma and arthritis). Both glinternet and pretrained lasso show improved performance in terms of ROC-AUC compared to the baselines trained on the ancestry-only and Mix datasets and produce the top performing ROC-AUC scores for all diseases except for cystitis and diabetes (Tables 4 and 5).
Both models perform particularly well for arthritis, with each combination of model and dataset giving a statistically significant p-value for the ROC tests. ROC-AUCs for those disease range from 0.062 (9.6 % increase) to 0.095 (15 % increase)
We find that L1-penalized Logistic Regression keeps 73 non-zero coefficients for arthritis, while pretrained lasso keeps only 2 of them, the PRS score and sex. It outperforms it by a ROC-AUC of 0.095. For glinternet for the same disease, we report a ROC-AUC of 0.091 and find 60 interaction terms (27 with age, 25 with the PRS score for diabetes and 8 with sex).
The pretrained lasso models trained on the Mix dataset have similar predictive ability compared to the baselines, while the best performing model in terms of average ROC-AUC score is pretrained lasso trained on the All dataset. The glinternet models perform similarly well in terms of average ROC-AUC, trained on either the Mix or All datasets. Asthma and osteoarthritis both have two significant p-values. Others combinations for different diseases show superior or on par performance with the baselines in terms of average ROC-AUC scores and p-values close to the threshold for statistical significance.
-1.70in0in Method Data CRF CYS GL DIA MI OST AST ART Avg. Inc. Logistic regression SA 0.748 0.5 0.5 0.782 0.675 0.706 0.5 0.706 0.639 0 Logistic regression Mix 0.834 0.663 0.629 0.802 0.703 0.688 0.624 0.612 0.694 0.055 Logistic regression All 0.83 0.555 0.635 0.815 0.714 0.691 0.639 0.646 0.691 0.051 Glinternet Mix 0.834 0.631 0.659 0.814 0.718 0.736 0.652 0.703 0.718 0.079 Glinternet All 0.84 0.629 0.648 0.813 0.725 0.739 0.649 0.708 0.719 0.079 Pretrained lasso Mix 0.829 0.63 0.551 0.807 0.746 0.714 0.642 0.706 0.703 0.064 Pretrained lasso All 0.84 0.634 0.648 0.811 0.75 0.713 0.645 0.726 0.721 0.081
-1.2in0in Method Metric Data CRF CYS GL DIA MI OST AST ART Glinternet P-value Mix 0.516 0.702 0.118 0.097 0.142 0.005 0.070 0.007 Glinternet AUC Mix -0.001 -0.032 0.03 0.012 0.015 0.048 0.028 0.091 Glinternet P-value All 0.157 0.074 0.221 0.607 0.210 0.015 0.273 0.020 Glinternet AUC All 0.01 0.075 0.013 -0.003 0.01 0.047 0.01 0.062 Pretrained lasso P-value Mix 0.747 0.708 0.772 0.334 0.113 0.226 0.039 0.023 Pretrained lasso AUC Mix -0.005 -0.033 -0.078 0.004 0.043 0.026 0.018 0.095 Pretrained lasso P-value All 0.200 0.064 0.391 0.746 0.129 0.168 0.187 0.016 Pretrained lasso AUC All 0.01 0.079 0.013 -0.005 0.036 0.022 0.006 0.08
Results for Admixed ancestry
For Admixed ancestry, we find only 2 statistically significant models (), both of which are pretrained lasso models for gall stones. We find that the Admixed glinternet and pretrained lasso models both perform better than the baselines but do not show statistically significant improvements (Tables 6 and 7). Pretrained lasso outperforms both baselines by a ROC-AUC ranging from 0.036 (4.8 % increase) for both datasets. Both glinterned and pretrained lasso are the top performing models in terms of average ROC-AUC except for myocardial infarction and chronic renal failure.
-1.75in0in Method Data CRF CYS GL DIA MI OST AST ART Avg. Inc. Logistic regression AD 0.883 0.5 0.664 0.902 0.816 0.681 0.641 0.5 0.698 0 Logistic regression Mix 0.91 0.666 0.668 0.882 0.774 0.682 0.638 0.651 0.734 0.036 Logistic regression All 0.909 0.666 0.675 0.892 0.777 0.7 0.645 0.672 0.742 0.044 Glinternet Mix 0.906 0.692 0.711 0.885 0.793 0.727 0.663 0.653 0.754 0.055 Glinternet All 0.892 0.681 0.667 0.884 0.795 0.754 0.673 0.693 0.755 0.056 Pretrained lasso Mix 0.905 0.685 0.84 0.948 0.752 0.7 0.656 0.67 0.77 0.071 Pretrained lasso All 0.906 0.65 0.866 0.947 0.757 0.719 0.666 0.706 0.777 0.079
-1.2in0in Method Metric Data CRF CYS GL DIA MI OST AST ART Glinternet P-value Mix 0.673 0.080 0.167 0.421 0.356 0.083 0.684 0.299 Glinternet AUC Mix -0.003 0.026 0.043 0.002 0.019 0.045 0.024 0.002 Glinternet P-value All 0.877 0.176 0.174 0.656 0.381 0.064 0.670 0.285 Glinternet AUC All -0.017 0.015 -0.008 -0.007 0.018 0.054 0.028 0.021 Pretrained lasso P-value Mix 0.630 0.175 0.001 0.092 0.570 0.320 0.787 0.269 Pretrained lasso AUC Mix -0.004 0.02 0.172 0.066 -0.023 0.018 0.018 0.018 Pretrained lasso P-value All 0.764 0.761 0.001 0.282 0.656 0.316 0.748 0.520 Pretrained lasso AUC All -0.003 -0.016 0.19 0.055 -0.02 0.018 0.02 0.034
Runtimes
Results of glinternet for osteoarthritis for South Asian ancestry
As seen on Figure 2, the PRS score is the main covariate from which interaction terms are found. The largest interaction term found is between the PRS score for osteoarthritis and the PC4. It has a coefficient of 26.07 and is three times larger than the second largest one (variables are standardized by glinternet when fitting). In a previous study, it was found that the 4th genetic PC of the UK Biobank helped to distinguish between the different South Asian groups (Indian, Pakistani, Bangladeshi and other South Asian and mixed backgrounds) [21]. This could mean different subgroups within the South Asian ancestry have varying predispositions to osteoarthritis.
-2.35in0in
Discussion
For 96 models trained, 16 had statistically significant incremental predictive performance in terms of ROC-AUC scores (), specifically for diabetes, arthritis, gall stones, cystitis, asthma and osteoarthritis. This could indicate a strong genetic and ancestral component for the prediction of those diseases. Our findings indicate that both interaction terms and pre-training can enhance prediction accuracy but for a limited set of diseases and moderate improvements in accuracy.
Pretrained lasso models were much more sparse than their L1-penalized logistic regression counterparts, making them easier to interpret and thus a superior candidate for prediction of certain diseases as they showed superior performance on average for South Asian and Admixed ancestries. They also performed better in terms of average ROC-AUC scores and p-values when trained on the All datasets versus the Mix ones.
We saw minor or even no significant improvements with the pretrained lasso models for African ancestry. This could indicate that people from the African ancestry exhibit distinct statistical characteristics compared to other subgroups, suggesting significant variations in the dataset and making it more difficult to generalize patterns from other ancestries.
Regarding the analysis of glinternet for osteoarthritis for South Asian ancestry, adding more interaction terms between the PRS score and other variables helped reach optimal performance in terms of cross-validation score, as main effects were still stable when relaxing the level of regularization. The main interaction coefficient driving the prediction was between the PRS score and the 4th PC, which was used in previous studies to distinguish betweeen different ethnicities within the South Asian ancestry.
Finally, both glinternet and pretrained lasso models took a significant amount of time to train compared to the L1-penalized logistic regression baselines. This can be explained by several factors. A pretrained lasso model is made of one glmnet model trained on all available ancestries and glmnet sub-models for each ancestry used. This can make training computationally expensive, as each new ancestry added requires a separate sub-model as well as increases the size of the training set for the overall model. For glinternet, as it evaluates pairwise variable interactions for the Group-LASSO model, the complexity scales up significantly when evaluating a large number of candidate interaction terms.
@@ Discussion
-1.05in0in
Fig 3: Diagram of the methodology followed for this project.
@@
Pretrained lasso
In Generalized Linear Models (GLMs) and -regularized GLMs, an offset—a predetermined -vector—can be added as an extra column in the feature matrix with a fixed weight of 1 and allows to control for a variable without estimating its parameter [22]. Additionally, the standard norm can be extended to a weighted norm by assigning a penalty factor to each feature, modulating the regularization applied [23]. A penalty factor of zero indicates no penalty, ensuring feature inclusion, while infinity results in feature exclusion, enhancing model flexibility and specificity.
Fit an overall LASSO model to the training data with all selected ancestries, selecting along the path that minimizes CV error. Set a fixed . Compute the offset and penalty factor for each input group as follows:
Calculate the linear predictor for each group , and define the offset as .
Identify the support set of . For each feature , define the penalty factor as .
For each ancestry , fit an ancestry-specific model.
Using the defined offset and penalty factor, train an L1-penalized logistic regression model on the data of the specific ancestry.
Use these models for group-specific predictions.
@@
The UK Biobank is a comprehensive cohort study that gathers data from various locations across the UK. To address the potential variability caused by population structure within our dataset, we limited our analysis to individuals who are not related, adhering to four specific criteria outlined in the UK Biobank’s sample quality control file, ukb_sqc_v2.txt. These criteria are: (1) inclusion in principal component analysis (as indicated in the used_in_pca_calculation column); (2) individuals not identified as outliers based on heterozygosity and missing data rates (found in the het_missing_outliers column); (3) absence of suspected anomalies in sex chromosome number (mentioned in the putative_sex_chromosome_aneuploidy column); and (4) having no more than ten likely third-degree relatives (as per the excess_relatives column).
We refined our study population by analyzing genetic and self-reported ancestry data, using the UK Biobank’s genotype principal components (PCs), self-reported ancestry data (UK Biobank Field ID 21000), and the in_white_British_ancestry_subset column from the sample QC file. Our focus was on individuals self-identifying as white British (337,129), non-British European (24,905), African (6,497), South Asian (7,831), and East Asian (1,704). The classification into five groups involved a two-stage process, initially using genotype PC loadings to apply specific criteria for each group: (1) white British based on PC1 and PC2 thresholds and inclusion in the white British ancestry subset; (2) non-British European with similar PC thresholds, identified as white but not white British; (3) African, with distinct PC thresholds, excluding identities as Asian, White, Mixed, or Other; (4) South Asian and (5) East Asian, each with unique PC thresholds and exclusions of identifying as Black, White, Mixed, or Other.
@@
Our cohort consisted of 80,810 White British (WB), 1,911 South Asians (SA), 1,499 Africans (AF), 6,783 categorized as Admixed (AD) and 5,910 Non-British Europeans(NBE). A detailed breakdown is available in S2 Fig.. We identified 8 common diseases to use for our study for which we had a minimum of 20 positive cases per ancestry: diabetes (DIA), asthma (AST), chronic renal failure (CRF), osteoarthritis (OST), myocardial infarction (MI), cystistis (CYS), gall stones (GS) and arthritis (ART).
-0.2in0in Ancestry Cases DIA MI AST GS OST ART CYS CRF WB Positive 5479 3234 10909 4103 7579 4969 2725 2862 WB Negative 75331 77576 69901 76707 73231 75841 78085 77948 WB Total 80810 80810 80810 80810 80810 80810 80810 80810 SA Positive 460 151 285 69 101 122 56 93 SA Negative 1451 1760 1626 1842 1810 1789 1855 1818 SA Total 1911 1911 1911 1911 1911 1911 1911 1911 AF Positive 232 30 201 36 86 103 24 83 AF Negative 1267 1469 1298 1463 1413 1396 1475 1416 AF Total 1499 1499 1499 1499 1499 1499 1499 1499 AD Positive 632 268 932 285 492 379 234 213 AD Negative 6151 6515 5851 6498 6291 6404 6549 6570 AD Total 6783 6783 6783 6783 6783 6783 6783 6783 NBE Positive 419 197 768 262 433 279 178 165 NBE Negative 5491 5713 5142 5648 5477 5631 5732 5745 NBE Total 5910 5910 5910 5910 5910 5910 5910 5910
@@
The dataset labeled “All” contains data from all ancestries available in the UK Biobank, including White British (WB) as well as ancestries South Asian, African, Admixed and Non-British European. The “Mix” dataset refers to a combination of White British data with data from one other specific ancestry at a time (e.g. White British and South Asian). This approach allows us to explore the predictive performance of our models when they are trained on data representing the majority population (White British) alongside data from a minority ancestry. The term “ancestry-only data” is used to denote datasets consisting exclusively of data from a single, specific ancestry—such as South Asian (SA) for instance. By employing these distinct dataset configurations, our study aims to assess whether training models on all ancestries is beneficial compared to training on White British data and data from a single ancestry.
@@
We started our analysis with a baseline model — an L1-penalized logistic regression — identified as the standard approach for this type of analysis in our literature review [2]. We used the glmnet package [24] in R [25] and trained it with the three distinct datasets aforementioned: ancestry-only, Mix, and All (see Figure 3).
Subsequently, we applied the glinternet package in R to train glinternet models [26], selecting age, sex, ancestry, and PRS score as interaction candidates (see Figure 3). Training of glinternet models is conducted on both the Mix and All datasets to assess glinternet’s ability to extrapolate patterns across ancestries facing data limitations.
Last, we adopted a two-stage process to train the pretrained lasso model. We used the ptLasso package in R [27] which trains a generic LASSO model using glmnet, from which is derived the vector. This vector is then used to construct an offset parameter, facilitating the training of ancestry-specific models. A comprehensive pan-ancestry model is subsequently developed by combining these steps (see Figure 3) and is trained like the glinternet models on both the Mix and All datasets.